# A Large Language Model Enhanced Sequential Recommender for Joint Video and Comment Recommendation

### v1

@@ Abstract.
In online video platforms, reading or writing comments on interesting videos has become an essential part of the video watching experience. However, existing video recommender systems mainly model users’ interaction behaviors with videos, lacking consideration of comments in user behavior modeling.
In this paper, we propose a novel recommendation approach called LSVCR by leveraging user interaction histories with both videos and comments, so as to jointly conduct personalized video and comment recommendation. Specifically, our approach consists of two key components, namely sequential recommendation (SR) model and supplemental large language model (LLM) recommender. The SR model serves as the primary recommendation backbone (retained in deployment) of our approach, allowing for efficient user preference modeling. Meanwhile, we leverage the LLM recommender as a supplemental component (discarded in deployment) to better capture underlying user preferences from heterogeneous interaction behaviors. In order to integrate the merits of the SR model and the supplemental LLM recommender, we design a two-stage training paradigm. The first stage is personalized preference alignment, which aims to align the preference representations from both components, thereby enhancing the semantics of the SR model. The second stage is recommendation-oriented fine-tuning, in which the alignment-enhanced SR model is fine-tuned according to specific objectives. Extensive experiments in both video and comment recommendation tasks demonstrate the effectiveness of LSVCR. Additionally, online A/B testing on the KuaiShou platform verifies the actual benefits brought by our approach. In particular, we achieve a significant overall gain of 4.13% in comment watch time.
@@
✉ Corresponding author.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Information systems Language models
@@ 1. Introduction
Nowadays, recommender systems play an essential role in alleviating information overload by providing personalized recommendation services with high-quality content resources. As a typical application scenario, online video platforms (e.g., YouTube, TikTok, KuaiShou [^1] https://www.kuaishou.com/) employ recommender systems to deliver customized videos to users, in which they primarily concentrate on capturing users’ video watching interests based on user-video interaction logs (Rendle et al., 2009; Hidasi et al., 2016; Kang and McAuley, 2018; Lin et al., 2022).
@@
With the growth of online video communities, the textual comments posted to the videos become increasingly crucial for enhancing the overall watching experience of users, providing complement or divergence information compared to the original video content. Our statistics show that more than 60% of users on the Kuaishou platform regularly view corresponding comments when watching the videos and express their interests through comments. In light of this, this paper aims to leverage both video and comment data to improve the recommendation quality and user engagement.
@@
Figure 1. A snapshot of one micro-video and its corresponding comment area on KuaiShou.
@@
To be more precise, we propose to study the joint tasks of video and comment recommendation in the specific scenario of KuaiShou, which is a popular Chinese micro-video platform (similar to TikTok). Figure 1 presents a snapshot of a micro-video and its associated comments on KuaiShou. When a user is watching a video and clicks the “comment” button, the comment page below the video will unfold, displaying a list of comments to the user. To develop a comprehensive solution, we integrate the modeling of videos and comments into a unified framework, considering the following three aspects:
  1. (1)
    While comments provide supplemental textual information for a video, their semantic modeling should be grounded in the video content. For example, in Figure 1, the keywords “sunset” and “waterfall” in the comment complement the video features and should be understood within the context of “fairyland” and “beautiful things”.
  2. (2)
    User’s interaction and feedback behaviors on comments indicate potential interest in watching the corresponding video. When a user actively interacts with the comments of a particular video, he/she is more likely to enjoy videos with similar content or comments.
  3. (3)
    In KuaiShou, a user could make multiple interactions with videos (e.g., like,collect, share) and corresponding comments (like, reply). Both interaction histories can capture users’ personalized video and comment preferences, which might suggest delivering high-quality video and customized comments instead of monotonous hot comments.
@@
Recently, large language models (LLMs) have demonstrated great potential in recommender systems (Cui et al., 2022; Wang et al., 2023; Zheng et al., 2023; Bao et al., 2023), due to their excellent semantic understanding and knowledge reasoning capabilities. In our task setting, we have abundant textual input such as video title and comment content, which naturally motivates us to leverage the impressive semantic modeling capability of LLM for developing our approach. Another potential advantage of using LLM in our setting is that the interaction histories of videos and comments essentially provide heterogeneous expressions of user behaviors, and we can leverage LLM to reason about the underlying user preferences from the two kinds of signals. However, the unaffordable computation costs of LLM restrict its efficient deployment as an online recommender in large-scale industrial systems (Zhang et al., 2023b; Hou et al., 2023; Xu et al., 2024). Considering this issue, we only adopt LLM in training as a supplement to enhance the semantics of our recommendation backbone.
@@
To this end, in this paper, we propose a novel framework that consolidates the benefits of LLM with the Sequential recommendation model for joint Video and Comment Recommendation, namely LSVCR. Our framework consists of two key components: conventional sequential recommendation (SR) model and supplemental LLM recommender. To begin with, we develop our recommendation backbone (retained in deployment) based on the SR model to achieve efficient user preference modeling. Furthermore, we introduce a supplemental LLM recommender (discarded in deployment) to enhance the preference semantics of the SR model. By verbalizing both types of user interaction histories in the format of textual instructions, we aim to leverage LLM’s semantic understanding and knowledge reasoning capabilities for enhanced preference modeling. In order to integrate the merits of the above two components, we design a training paradigm consisting of two stages: personalized preference alignment and recommendation-oriented fine-tuning. For the alignment stage, we jointly learn the video and comment recommendation tasks while aligning the preference representation of the SR model with that of the supplemental LLM recommender. For the fine-tuning stage, we discard the supplemental LLM recommender and further adapt the alignment-enhanced SR model to specific tasks for better recommendation performance.
@@
To evaluate our approach, we conduct extensive experiments on a large real-world industrial dataset. The experimental results demonstrate the significant effectiveness of LSVCR compared to competitive baselines for both video and comment recommendation tasks. Additionally, online A/B testing verifies the benefits of LSVCR in the industrial recommendation system of KuaiShou. Especially in the comment end, we achieve a 4.13% increase in watch time and a 1.36% gain in interaction number.
@@ 2. Problem Formulation
In our task, we consider the online micro-video scenario, where users can freely watch interesting videos and further view comments below specific videos. Different from previous video recommendation studies (Rendle et al., 2009; Hidasi et al., 2016; Kang and McAuley, 2018; Lin et al., 2022), we consider both video recommendation and comment recommendation, in which comment recommendation is helpful to enhance the user engagement and improve the user experience. Formally, we denote the sets of users and videos as $\mathcal{U}$ and $\mathcal{V}$ respectively. Each video $v\in\mathcal{V}$ is associated with a textual title and a list of corresponding comments, denoted as $\mathcal{C}_{v}$. For a user $u\in\mathcal{U}$, there exist two chronological interaction sequences: (1) Video interactions $\mathcal{S}_{v}=[v_{1},v_{2},\dots,v_{n}]$, where $v_{i}$ is the $i$-th watched video by the user. (2) Comment interactions $\mathcal{S}_{c}=[r_{1},r_{2},\dots,r_{m}]$, where $r_{j}$ is the $j$-th comment interaction record of the user, and each record contains a set of comments that the user has interacted with below video $\tilde{v}_{j}$, denoted as $(c_{j}^{1},\dots,c_{j}^{k})$ In addition, we denote the original videos of these comments as $\mathcal{S}_{\tilde{v}}=[\tilde{v}_{1},\tilde{v}_{2},\dots,\tilde{v}_{m}]$ for further elaboration. In KuaiShou, a user can only engage in the comment thread without interacting with the corresponding video. Thus, the videos in $\mathcal{S}_{v}$ and $\mathcal{S}_{\tilde{v}}$ may not be the same, but likely share overlapped ones.
@@ 3. Methodology
In this section, we present LSVCR for joint video and comment recommendation. Its overall framework is depicted in Figure 2. We first introduce the SR model for sequential user preference modeling in Section 3.1. Then we incorporate the supplemental LLM recommender in Section 3.2 to enhance preference modeling. After that, a personalized preference alignment approach is proposed in Section 3.3 to align the SR model with the supplemental LLM recommender. Finally, we fine-tune the alignment-enhanced SR model for video or comment recommendation tasks in Section 3.4.
@@ 3.1.1. Text Feature Embedding with LLM
Following our problem formulation in Section 2, we use the user’s historical interaction sequences, i.e., $\mathcal{S}_{v}$ and $\mathcal{S}_{c}$, to construct inputs. First, we employ LLM to encode the titles and comments involved in the inputs. Then, we combine the output embeddings as the representation of each interaction, which can be formally written as:
@@
where $t_{i}$ and $c_{i}$ denote the title and a popular comment of video $v_{i}$ in $\mathcal{S}_{v}$ respectively. $t_{j}$ denotes the title of video $\tilde{v}_{j}$ corresponding to the interaction record $r_{j}$ in $\mathcal{S}_{c}$. “——” denotes concatenation operation, $\mathbf{W}_{1}\in\mathbb{R}^{2d\times d}$ is a trainable parameter for linear projection, and $\operatorname{MEAN}(\cdot)$ denotes mean pooling. $\mathbf{z}^{v}_{i}$ and $\mathbf{z}^{c}_{j}$ represent the text feature embeddings for one video/comment interaction, and the text feature embeddings of interaction sequences are denoted as $\mathbf{Z}^{v}\in\mathbb{R}^{n\times d}$ and $\mathbf{Z}^{c}\in\mathbb{R}^{m\times d}$.
@@ 3.1.2. Randomized Positional Encoding for Length Extension
In order to model the sequential pattern of user interaction histories, we incorporate two learnable positional encoding matrices, denoted as $\mathbf{P}^{v}\in\mathbb{R}^{L_{v}\times d}$ and $\mathbf{P}^{c}\in\mathbb{R}^{L_{c}\times d}$, respectively. Here $L_{v}$ and $L_{c}$ represent the maximum lengths of the historical interaction sequences. In conventional sequential recommendation, the length can generally be set longer (Hidasi et al., 2016; Kang and McAuley, 2018; Sun et al., 2019). However, in the personalized preference alignment training stage, it is set shorter which is limited by LLM input length and computational consumption (Geng et al., 2022; Zhang et al., 2023b; Zheng et al., 2023). Inspired by the length extension technique in the NLP field (Ruoss et al., 2023), we randomly select an ordered subset of positional encodings during the alignment stage, while applying the entire positional encodings in subsequent fine-tuning and inference processes. Formally, the positional encodings of $\mathcal{S}_{v}^{u}$ are sampled as follows:
@@ 3.1.3. Sequence Representation Learning
Given the user interaction histories and sampled position encodings, we use the Transformer backbone (Vaswani et al., 2017) to model video and comment interaction sequences. The above sequence embeddings (i.e., $\mathbf{E}^{v}$ and $\mathbf{E}^{c}$) and positional encodings (i.e., $\widetilde{\mathbf{P}}^{v}$ and $\widetilde{\mathbf{P}}^{c}$) are summed up as inputs of Transformer layers. The calculations can be formally written as:
(4) $\mathbf{H}^{v}$ $=\operatorname{Transformer}^{v}(\mathbf{E}^{v}+\widetilde{\mathbf{P}}^{v}),$
(5) $\mathbf{H}^{c}$ $=\operatorname{Transformer}^{c}(\mathbf{E}^{c}+\widetilde{\mathbf{P}}^{c}),$

### v2

# Enhancing Sequential Recommender with Large Language Models for Joint Video and Comment Recommendation

@@ Abstract.
Nowadays, reading or writing comments on captivating videos has emerged as a critical part of the viewing experience on online video platforms. However, existing recommender systems primarily focus on users’ interaction behaviors with videos, neglecting comment content and interaction in user preference modeling. In this paper, we propose a novel recommendation approach called LSVCR that utilizes user interaction histories with both videos and comments to jointly perform personalized video and comment recommendation. Specifically, our approach comprises two key components: sequential recommendation (SR) model and supplemental large language model (LLM) recommender. The SR model functions as the primary recommendation backbone (retained in deployment) of our method for efficient user preference modeling. Concurrently, we employ a LLM as the supplemental recommender (discarded in deployment) to better capture underlying user preferences derived from heterogeneous interaction behaviors. In order to integrate the strengths of the SR model and the supplemental LLM recommender, we introduce a two-stage training paradigm. The first stage, personalized preference alignment, aims to align the preference representations from both components, thereby enhancing the semantics of the SR model. The second stage, recommendation-oriented fine-tuning, involves fine-tuning the alignment-enhanced SR model according to specific objectives. Extensive experiments in both video and comment recommendation tasks demonstrate the effectiveness of LSVCR. Moreover, online A/B testing on KuaiShou platform verifies the practical benefits of our approach. In particular, we attain a cumulative gain of 4.13% in comment watch time.
@@
$\dagger$  Corresponding author.
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the Nineteenth ACM Conference on Recommender Systems; September 22–26, 2025; Prague, Czech Republicbooktitle: Proceedings of the Nineteenth ACM Conference on Recommender Systems (RecSys ’25), September 22–26, 2025, Prague, Czech Republicdoi: 10.1145/3705328.3748075isbn: 979-8-4007-1364-4/2025/09ccs: Information systems Recommender systems
@@ 1. Introduction
Nowadays, recommender systems play an essential role in alleviating information overload by providing personalized recommendation services with high-quality content resources. As a typical application scenario, online video platforms (e.g., YouTube, TikTok, KuaiShou) employ recommender systems to deliver customized videos to users, in which they primarily concentrate on capturing users’ video watching interests based on user-video interaction logs (Rendle et al., 2009; Hidasi et al., 2016; Kang and McAuley, 2018; Lin et al., 2022). With the growth of online video communities, the textual comments posted to the videos become increasingly crucial for enhancing the overall watching experience of users, providing complement or divergence information compared to the original video content. Our statistics show that more than 60% of users on our video platform regularly view corresponding comments when watching the videos and express their interests through comments. In light of this, this paper aims to leverage both video and comment data to improve the recommendation quality and user engagement.
@@
Figure 1. A snapshot of one micro-video and its corresponding comment area on KuaiShou platform.
@@
To be more precise, we propose to study the joint tasks of video and comment recommendation in the specific scenario of a micro-video platform (e.g., KuaiShou, TikTok ). Figure 1 presents a snapshot of a micro-video and its associated comments on our platform. When a user is watching a video and clicks the “comment” button, the comment page below the video will unfold, displaying a list of comments to the user. To develop a comprehensive solution, we integrate the modeling of videos and comments into a unified framework, considering the following three aspects: (1) While comments provide supplemental textual information for a video, their semantic modeling should be grounded in the video content. For example, in Figure 1, the keywords “sunset” and “waterfall” in the comment complement the video features and should be understood within the context of “fairyland” and “beautiful things”. (2) User’s interactions and feedback on comments indicate potential interest in watching the corresponding video. When a user actively interacts with the comments of a particular video, he/she is more likely to enjoy videos with similar content or comments. (3) In our video platform, a user could make multiple interactions with videos (e.g., like,collect, share) and corresponding comments (like, reply). Both interaction histories can capture users’ personalized video and comment preferences, which might suggest delivering high-quality video and customized comments instead of monotonous hot comments.
@@
Recently, large language models (LLMs) have demonstrated great potential in recommender systems (Cui et al., 2022; Wang et al., 2023; Zheng et al., 2023; Bao et al., 2023), due to their excellent semantic understanding and knowledge reasoning capabilities. In our task setting, we have abundant textual input such as video title and comment content, which naturally motivates us to leverage the impressive semantic modeling capability of LLMs for developing our approach. Another potential advantage of using LLMs in our setting is that the interaction histories of videos and comments essentially provide heterogeneous expressions of user behaviors. Thus, we can leverage LLMs to reason about the underlying user preferences from the two kinds of interactions. However, the unaffordable computation costs of LLMs restrict their deployment as the online recommender in large-scale industrial systems (Zhang et al., 2023b; Hou et al., 2023b; Xu et al., 2024). Considering this issue, we only adopt a LLM in training as a supplement to enhance the semantics of our recommendation backbone.
@@
To this end, in this paper, we propose a novel framework that consolidates the benefits of LLM with the Sequential recommendation model for joint Video and Comment Recommendation, namely LSVCR. Our framework consists of two key components: conventional sequential recommendation (SR) model and supplemental LLM recommender. To begin with, we develop our recommendation backbone (retained in deployment) based on the SR model to achieve efficient user preference modeling. Furthermore, we introduce a supplemental LLM recommender (discarded in deployment) to enhance the preference semantics of the SR model. By verbalizing both types of user interaction histories in the format of textual instructions, we aim to leverage the LLM’s semantic understanding and knowledge reasoning capabilities for enhanced preference modeling. In order to integrate the merits of the above two components, we design a training paradigm consisting of two stages: personalized preference alignment and recommendation-oriented fine-tuning. For the alignment stage, we jointly learn the video and comment recommendation tasks while aligning the preference representation of the SR model with that of the supplemental LLM recommender. For the fine-tuning stage, we discard the supplemental LLM recommender and further adapt the alignment-enhanced SR model to specific tasks for better recommendation performance.
@@
To evaluate our approach, we conduct extensive experiments on a large real-world industrial dataset and a public Amazon benchmark. The experimental results demonstrate the significant effectiveness of LSVCR compared to competitive baselines for both video and comment recommendation tasks. Additionally, online A/B testing verifies the benefits of LSVCR in the industrial recommender system. Especially in the comment end, we achieve a 4.13% increase in watch time and a 1.36% gain in interaction number.
2. Related Work
Sequential Recommendation. Sequential recommendation aims to mine users’ personalized preferences and recommend the next item for each user. This field has attracted a lot of studies due to its advantages in capturing the dynamic characteristics of user behaviors (Hidasi et al., 2016; Kang and McAuley, 2018; Sun et al., 2019). Early methods follow the Markov Chain assumption and apply matrix factorization to model item-item transfer relationships (Rendle et al., 2010; He and McAuley, 2016). Recently, typical works are mostly based on various deep neural networks to better user behavior modeling, including RNN (Hidasi et al., 2016; Li et al., 2017), CNN (Tang and Wang, 2018; Yuan et al., 2019), GNN (Wu et al., 2019a; Xu et al., 2019), and Transformer (Kang and McAuley, 2018; Sun et al., 2019). Furthermore, several studies focus on utilizing additional contextual information (e.g., title, description, category) to enhance item sequence modeling (Zhang et al., 2019; Zhou et al., 2020; Xie et al., 2022). Unlike these works, this paper specifically concentrates on video comments, incorporating abundant textual information and user feedback from comment pages into sequential recommendation.
Text-Enhanced Recommendation. On various application platforms, items are associated with abundant textual features, which motivates researchers to engage in text-enhanced recommendation models. Recently, large language models (LLMs) have gained significant popularity and shown great potential in wide-ranging fields (Zhao et al., 2023; Jiang et al., 2023; Wu et al., 2023). In the realm of text-enhanced recommendation, LLM generally plays distinct roles: (1) LLM serves as a text feature encoder, allowing the text features to be encoded asynchronously. Subsequently, a conventional recommendation model utilizes knowledge-aware text embeddings for recommendation (Wu et al., 2021; Hou et al., 2022; Yuan et al., 2023; Harte et al., 2023; Ren et al., 2023; Hou et al., 2023a). Although effective, this approach often requires special mechanisms to enhance text feature modeling for better performance, such as contrastive learning (Hou et al., 2022; Ren et al., 2023). (2) LLM serves as a recommender. This approach typically constructs the user’s interaction history and other contexts into textual instructions, and then LLM is instructed to directly generate recommendation targets (Zhang et al., 2023b; Hou et al., 2023b; Xu et al., 2024; Zheng et al., 2023; Bao et al., 2023). More recently, several works leverage various methods to improve the adaptation of LLM for recommendation scenarios, including incorporating collaborative embedding (Zhang et al., 2023a; Li et al., 2023; Liao et al., 2023), and semantic alignment (Zhu et al., 2023; Zheng et al., 2023). These methods fully utilize the excellent capabilities of LLM and achieve astonishing results, but they still cannot avoid the high computational cost and slow inference time caused by LLM. As a comparison, we aim to consolidate the merits of these two approaches through personalized preference alignment in order to achieve effective textual semantics understanding and heterogeneous interaction behavior modeling.
@@ 3. Problem Formulation
In our task, we consider the online micro-video scenario, where users can freely watch interesting videos and further view comments below specific videos. Different from previous video recommendation studies (Rendle et al., 2009; Hidasi et al., 2016; Kang and McAuley, 2018; Lin et al., 2022), we consider both video recommendation and comment recommendation, in which comment recommendation is helpful to enhance the user engagement and improve the user experience. Formally, we denote the sets of users and videos as $\mathcal{U}$ and $\mathcal{V}$ respectively. Each video $v\in\mathcal{V}$ is associated with a textual title and a list of corresponding comments, denoted as $\mathcal{C}_{v}$. For a user $u\in\mathcal{U}$, there exist two chronological interaction sequences: (1) Video interactions $\mathcal{S}_{v}=[v_{1},v_{2},\dots,v_{n}]$, where $v_{i}$ is the $i$-th watched video by the user. (2) Comment interactions $\mathcal{S}_{c}=[r_{1},r_{2},\dots,r_{m}]$, where $r_{j}$ is the $j$-th comment interaction record of the user, and each record contains a set of comments that the user has interacted with below video $\tilde{v}_{j}$, denoted as $(c_{j}^{1},\dots,c_{j}^{k})$ In addition, we denote the original videos of these comments as $\mathcal{S}_{\tilde{v}}=[\tilde{v}_{1},\tilde{v}_{2},\dots,\tilde{v}_{m}]$ for further elaboration. In our platform, a user can only engage in the comment thread without interacting with the corresponding video. Thus, the videos in $\mathcal{S}_{v}$ and $\mathcal{S}_{\tilde{v}}$ may not be the same, but likely share overlapped ones.
@@ 4. Methodology
In this section, we present our proposed LSVCR for joint video and comment recommendation. Its overall framework is depicted in Figure 2. We first introduce the SR model for sequential user preference modeling in Section 4.1. Then we incorporate the supplemental LLM recommender in Section 4.2 to enhance preference modeling. After that, a personalized preference alignment approach is proposed in Section 4.3 to align the SR model with the supplemental LLM recommender. Finally, we fine-tune the alignment-enhanced SR model for video or comment recommendation tasks in Section 4.4.
@@ 4.1.1. Text Feature Embedding with LLM
Following our problem formulation, we use the user’s interaction sequences, i.e., $\mathcal{S}_{v}$ and $\mathcal{S}_{c}$, to construct inputs. First, we employ LLM to encode the titles and comments involved in the inputs. Then, we combine the outputs as the representation of each interaction, which can be written as:
@@
where $t_{i}$ and $c_{i}$ denote the title and a hot comment of video $v_{i}$ in $\mathcal{S}_{v}$. $t_{j}$ denotes the title of video $\tilde{v}_{j}$ corresponding to the interaction record $r_{j}$ in $\mathcal{S}_{c}$. “——” denotes concatenation operation, $\mathbf{W}_{1}\in\mathbb{R}^{2d\times d}$ is a trainable parameter for linear projection, and $\operatorname{MEAN}(\cdot)$ denotes mean pooling. $\mathbf{z}^{v}_{i}$ and $\mathbf{z}^{c}_{j}$ represent the text feature embeddings for one video/comment interaction, and the text feature embeddings of interaction sequences are denoted as $\mathbf{Z}^{v}\in\mathbb{R}^{n\times d}$ and $\mathbf{Z}^{c}\in\mathbb{R}^{m\times d}$.
@@ 4.1.2. Randomized Positional Encoding for Length Extension
To model the sequential pattern of user histories, we maintain two learnable positional encoding matrices, denoted as $\mathbf{P}^{v}\in\mathbb{R}^{L_{v}\times d}$ and $\mathbf{P}^{c}\in\mathbb{R}^{L_{c}\times d}$, respectively. $L_{v}$ and $L_{c}$ are the maximum lengths of the interaction sequences. In conventional sequential recommendation, the length can generally be set longer (Hidasi et al., 2016; Kang and McAuley, 2018). However, in the personalized preference alignment stage, it is set shorter, limited by LLM input length and computational consumption. Inspired by the length extension technique in the NLP (Ruoss et al., 2023), we randomly select an ordered subset of positional encodings during the alignment stage, while applying the entire set in fine-tuning and inference processes. Formally, the positional encodings of $\mathcal{S}_{v}^{u}$ are sampled as follows:
@@ 4.1.3. Sequence Representation Learning
Given the user interaction histories and sampled position encodings, we use the Transformer backbone (Vaswani et al., 2017) to model video and comment interaction sequences. The above sequence embeddings (i.e., $\mathbf{Z}^{v}$ and $\mathbf{Z}^{c}$) and positional encodings (i.e., $\widetilde{\mathbf{P}}^{v}$ and $\widetilde{\mathbf{P}}^{c}$) are summed up as inputs of Transformer layers. The calculations can be formally written as:
(4) $\mathbf{H}^{v}=\operatorname{TRM}^{v}(\mathbf{Z}^{v}+\widetilde{\mathbf{P}}^{v}),\ \ \ \ \mathbf{H}^{c}=\operatorname{TRM}^{c}(\mathbf{Z}^{c}+\widetilde{\mathbf{P}}^{c}),$